# Deflated Dynamics Value Iteration

### v1

@@ 10.  Environments
Figure 5: Chain walk environment. Dark Shaded region is goal state with a positive reward and grey shaded region is state with a negative reward. Arrows show optimal policy.
Figure 6: Cliffwalk environment. Dark shaded region is goal state with a positive reward. Gray shaded regions are terminal states with negative reward. Arrows show optimal policy.
@@
The Cliffwalk is a $3\times 7$ grid world. The top-left corner is initial state, top-right corner is goal state with reward of $10$, and other states in first row are terminal states with reward of $-10$. There is a penalty of $-1$ in all other states. The agent has four actions: UP(00), RIGHT($1$), DOWN($2$), and LEFT($3$). Each action has $90\%$ chance to successfully move, but with probability of $10\%$ one of the other three directions is randomly chosen and the agent moves in that direction. If the agent attempts to get out of the boundary, it will stay in place. Optimal and non optimal poilcies of Cliffwalk used in experiments are as follows:
Optimal policy: $[2,0,0,0,0,0,0,2,2,2,2,2,1,0,1,1,1,1,1,1,0]$
Non optimal policy: $[0,0,2,1,2,1,2,1,0,2,0,3,2,0,3,1,0,1,1,1,3]$
@@
The Maze is a $5\times 5$ grid world with $16$ walls. The top-left corner is the initial state, and the bottom-left corner is the goal state with reward of $10$. Similar to the Clifffwalk, there is a penalty of $-1$ in all other states, and the agent has four actions: UP(00), RIGHT($1$), DOWN($2$), and LEFT($3$). Each action has $90\%$ chance to successfully move but with probability of $10\%$ one of the other three directions is randomly chosen and the agent moves in that direction. If the agent attempts to get out of the boundary or hits a wall, it will stay in place. Optimal and non optimal poilcies of Maze used in experiments are as follows:
Optimal policy: $[2,1,1,1,2,2,0,2,3,2,2,0,2,0,3,1,0,1,1,2,0,3,3,3,3]$
Non optimal policy: $[2,2,3,0,3,0,2,1,3,2,2,2,3,3,1,0,3,0,3,3,3,3,1,1,0]$
@@
We use the Chain Walk environment, as described by Farahmand and Ghavamzadeh [2021], which is similar to the formulation by Lagoudakis and Parr [2003]. Chain Walk is parametrized by the tuple $(|\mathcal{X}|,|\mathcal{A}|,b_{p},b_{r})$. It is circular chain, where the state $1$ and $50$ are connected. The reward is state-dependent. State $3$ is goal state with reward $1$, state $50$ gives reward $-1$, and all other states give reward 00. Agent has two actions: RIGHT(00) and LEFT($1$). Each action has $80\%$ chance to successfully move but with probability of $40/3\%$ agent will stay in place, and with probability of $20/3\%$ agent will move opposite direction. Optimal poilcy of Chain Walk used in experiments is as follows:
Optimal policy: [0,0,0,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0,0,0,0,0,[0,0,0,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0,0,0,0,0,[ 0 , 0 , 0 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 0 , 0 , 0 , 0 , 0 , 0 , 0 , 0 , 0 , 0,0,0,0,0,0,0,0,0,0,0,0]0,0,0,0,0,0,0,0,0,0,0,0]0 , 0 , 0 , 0 , 0 , 0 , 0 , 0 , 0 , 0 , 0 , 0 ]
Non optimal policy: [1,0,0,1,1,1,1,1,1,1,0,1,0,0,0,0,0,0,1,1,1,1,1,1,0,1,1,1,0,0,0,1,1,1,1,1,[1,0,0,1,1,1,1,1,1,1,0,1,0,0,0,0,0,0,1,1,1,1,1,1,0,1,1,1,0,0,0,1,1,1,1,1,[ 1 , 0 , 0 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 0 , 1 , 0 , 0 , 0 , 0 , 0 , 0 , 1 , 1 , 1 , 1 , 1 , 1 , 0 , 1 , 1 , 1 , 0 , 0 , 0 , 1 , 1 , 1 , 1 , 1 ,
0,0,1,1,0,1,0,0,1,1,0,0,1,0]0,0,1,1,0,1,0,0,1,1,0,0,1,0]0 , 0 , 1 , 1 , 0 , 1 , 0 , 0 , 1 , 1 , 0 , 0 , 1 , 0 ], [0,1,1,0,1,1,0,0,1,1,0,1,1,1,1,0,1,1,1,0,0,0,0,0,1,1,1,1,1,1,1,[0,1,1,0,1,1,0,0,1,1,0,1,1,1,1,0,1,1,1,0,0,0,0,0,1,1,1,1,1,1,1,[ 0 , 1 , 1 , 0 , 1 , 1 , 0 , 0 , 1 , 1 , 0 , 1 , 1 , 1 , 1 , 0 , 1 , 1 , 1 , 0 , 0 , 0 , 0 , 0 , 1 , 1 , 1 , 1 , 1 , 1 , 1 ,
1,1,1,1,1,1,0,0,1,0,0,1,0,1,0,1,0,1,1]1,1,1,1,1,1,0,0,1,0,0,1,0,1,0,1,0,1,1]1 , 1 , 1 , 1 , 1 , 1 , 0 , 0 , 1 , 0 , 0 , 1 , 0 , 1 , 0 , 1 , 0 , 1 , 1 ]
Garnet
We use the Garnet environment as described by Farahmand and Ghavamzadeh [2021], Rakhsha et al. [2022], which is based on Bhatnagar et al. [2009]. Garnet is parameterized by the tuple $(|\mathcal{X}|,|\mathcal{A}|,b_{p},b_{r})$. $|\mathcal{X}|$ and $|\mathcal{A}|$ are the number of states and actions, and $b_{p}$ is the branching factor of the environment, which is the number of possible next states for each state-action pair. We randomly select $b_{p}$ states without replacement and then, transition distribution is generated randomly. We select $b_{r}$ states without replacement, and for each selected $x$, we assign state-dependent reward a uniformly sampled value in $(0,1)$.
11.  Rank-$1$ DDVI for Control Experiments
@@
12.  Additional DDVI Experiments and Details
In all experiments, we set DDVI’s $\alpha=0.99$. In experiments for Figure 1, the QR iteration is run for 600 iterations. For PID VI, we set $\eta=0.05$ and $\epsilon=10^{-10}$. In Anderson VI, we have $m=5$. In Figure 2, we use 20 Garnet mdps with branching factor $b_{p}=3$, and $b_{r}=0.1*|\mathcal{X}|$.
We perform further comparison of convergence in Figures 8, 9, 10, 11. Figure 8 is run with Garnet environment with $50$ states and $40$ branching factor that matches the setting in [Goyal and Grand-Clément, 2022].
@@ 12.  Additional DDVI Experiments and Details
Figure 10: Comparison of DDVI with other accelerated VIs. Normalized errors are shown against the iteration number (left) and wall clock time (right) in Chain Walkk.
@@
13.  DDTD Experiments
A key part of our experimental setup is the model $\hat{\mathcal{P}}$. To show the robustness of DDTD to model error and also its advantage over Dyna, we consider the scenario that $\hat{\mathcal{P}}$ has some non-diminishing error. We achieve this with the same technique as [Rakhsha et al., 2022]. Assume that each iteration $t$, the empirical distribution of the next state from $x,a$ is $\mathcal{P}_{\mathrm{MLE}}(\cdot|x,a)$, which is going to be updated with every sample. For some hyperparameter $\theta\in[0,1]$, we set the approximate dynamics $\hat{\mathcal{P}}(\cdot|x,a)$ as
@@ 13.  DDTD Experiments
As we see in Figure 12 and Figure 13, DDTD shows a faster convergence than the conventional TD. In Maze, this is only achieved with higher rank versions of DDTD but in Chain Walk, even rank-1 DDTD is able to significantly accelerate the learning. Also note that unlike Dyna, DDTD is converging to the true value function despite the model error. We observe that the impact of model error on DDTD is very mild. In some cases such as rank-3 DDTD in Maze, between $\theta=0.3$ and $\theta=0.5$, we observe slightly slower convergence with higher model error. The hyperparamaters of TD Learning and DDTD are given in Table 1 and 2.
Figure 12: Comparison of DDTD with TD Learning and Dyna in Maze. Left: low model error with $\theta=0.1$. Middle: medium model error with $\theta=0.3$. Right: high model error with $\theta=0.5$. Each curve is average of 20 runs. Shaded area shows one standard error.
Table 1: Hyperparamters for the Maze environment. Cells with multiple values provide the value of the hyperparameter for $\theta=0.1$, $\theta=0.3$, and $\theta=0.5$, respectively.
@@
Figure 13: Comparison of DDTD with TD Learning and Dyna in Chainwalk. Left: low model error with $\theta=0.1$. Middle: medium model error with $\theta=0.3$. Right: high model error with $\theta=0.5$. Each curve is average of 20 runs. Shaded area shows one standard error.
Table 2: Hyperparamters for the Chainwalk environment. Cells with multiple values provide the value of the hyperparameter for $\theta=0.1$, $\theta=0.3$, and $\theta=0.5$, respectively.

### v2

@@ Appendix D Environments
Figure 6: Chain walk environment. Dark shaded region is the goal state with a positive reward and gray shaded region is state with a negative reward. Arrows show optimal policy.
@@
The Cliffwalk is a $3\times 7$ grid world. The top-left corner is initial state, top-right corner is the goal terminal state with reward of $10$, and the other states in the first row are terminal states with a reward of $-10$. There is a penalty of $-1$ in all other states. The agent has four actions: UP(00), RIGHT($1$), DOWN($2$), and LEFT($3$). Each action has $90\%$ chance to successfully move in the selected direction, and with probability of $10\%$ one of the other three directions is randomly selected and the agent moves in that direction. If the agent attempts to get out of the boundary, it will stay in place. We use the optimal policy for the policy evaluation task in our experiments. The environment is shown in Figure 4.
@@
The Maze is a $5\times 5$ grid world with $16$ walls. The top-left corner is the initial state, and the bottom-left corner is the goal state with reward of $10$. Similar to the Clifffwalk, there is a penalty of $-1$ in all other states, and the agent has four actions: UP(00), RIGHT($1$), DOWN($2$), and LEFT($3$). Each action has $90\%$ chance to successfully move in the selected direction, and with probability of $10\%$ one of the other three directions is randomly selected and the agent moves in that direction. If the agent attempts to get out of the boundary or hits a wall, it will stay in place. We use the following policies for the policy evaluation task in our DDVI and DDTD experiments, respectively:
$(2,2,3,0,3,0,2,1,3,2,2,2,3,3,1,0,3,0,3,3,2,2,1,1,0),$
$(2,2,3,0,3,0,2,1,3,2,2,2,3,3,1,0,3,0,3,3,3,3,1,1,0).$
Here, the action for each state is listed row-wise. The environment is shown in Figure 5.
@@
We use the Chain Walk environment, as described by Farahmand & Ghavamzadeh (2021), which is similar to the formulation by Lagoudakis & Parr (2003). Chain Walk is parametrized by the tuple $(|\mathcal{X}|,|\mathcal{A}|,b_{p},b_{r})$. It is a circular chain, where the state $1$ and $50$ are connected. The reward is state-dependent. State $40$ is goal state with reward $1$, state $11$ gives reward $-1$, and all other states give reward 00. Agent has two actions: RIGHT(00) and LEFT($1$). With each action, the agent has $70\%$ chance to successfully move in the selected direction, $10\%$ chance to stay still, and with probability of $20\%$ the agent moves in the opposite direction. We use the following policy in our experiments: (0, 1, 1, 0, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 0, 1, 0, 1, 1). The environment is shown in Figure 6.
@@
Garnet
We use the Garnet environment as described by Farahmand & Ghavamzadeh (2021); Rakhsha et al. (2022), which is based on Bhatnagar et al. (2009). Garnet is parameterized by the tuple $(|\mathcal{X}|,|\mathcal{A}|,b_{p},b_{r})$. $|\mathcal{X}|$ and $|\mathcal{A}|$ are the number of states and actions, and $b_{p}$ is the branching factor of the environment, which is the number of possible next-states for each state-action pair. We randomly select $b_{p}$ states without replacement and then, transition distribution is generated randomly. We select $b_{r}$ states without replacement, and for each selected $x$, we assign state-dependent reward to be a uniformly sampled value in $(0,1)$.
Appendix E Rank-$1$ DDVI for Control Experiments
@@
Appendix F Additional DDVI Experiments and Details
In all experiments, we set DDVI’s $\alpha=0.99$. We perform further experiments of DDVI with different ranks for discount factor $\gamma=0.95,0.995$ (Figure 8 and  9), and the QR iteration is run for 100 iterations in experiments.
Figure 8: Comparison of DDVI with different ranks, AutoPI, and AutoQR against VI with discount factor $\gamma=0.95$ in (left) Maze and (right) Chain Walk. The plots for DDVI do not start at iteration 00 because the costs of computing $E_{s}$ through QR iterations are incorporated as rightward shifts. Rate of DDVI with AutoPI and AutoQR changes when $E_{s}$ is updated.
Figure 9: Comparison of DDVI with different ranks, AutoPI, and AutoQR against VI with discount factor $\gamma=0.995$ in (left) Maze and (right) Chain Walk. The plots for DDVI do not start at iteration 00 because the costs of computing $E_{s}$ through QR iterations are incorporated as rightward shifts. Rate of DDVI with AutoPI and AutoQR changes when $E_{s}$ is updated.
@@ Appendix F Additional DDVI Experiments and Details
Figure 12: Comparison of DDVI with other accelerated VIs. Normalized errors are shown against the iteration number (left) and wall clock time (right) in Chain Walk.
@@
We perform further comparison of convergence in Figures 10, 11, 12, 13. Figure 10 is run with Garnet environment with $50$ states and $40$ branching factor that matches the setting in Goyal & Grand-Clément (2022). In experiments, the QR iteration is run for 600 iterations. For PID VI, we set $\eta=0.05$ and $\epsilon=10^{-10}$. In Anderson VI, we have $m=5$. In Figure 2, we use 20 Garnet MDPs with branching factor $b_{p}=2$, and $b_{r}=0.1|\mathcal{X}|$.
Appendix G DDTD Experiments
A key part of our experimental setup is the model $\hat{\mathcal{P}}$. To show the robustness of DDTD to model error and also its advantage over Dyna, we consider the scenario that $\hat{\mathcal{P}}$ has some non-diminishing error. We achieve this with the same technique as Rakhsha et al. (2022). Assume that each iteration $t$, the empirical distribution of the next- state from $x,a$ is $\mathcal{P}_{\mathrm{MLE}}(\cdot|x,a)$, which is going to be updated with every sample. For some hyperparameter $\theta\in[0,1]$, we set the approximate dynamics $\hat{\mathcal{P}}(\cdot|x,a)$ as
@@ Appendix G DDTD Experiments
As we see in Figure 14 and Figure 15, DDTD shows a faster convergence than the conventional TD. In Maze, this is only achieved with higher rank versions of DDTD but in Chain Walk, even rank-1 DDTD is able to significantly accelerate the learning. Also note that unlike Dyna, DDTD is converging to the true value function despite the model error. We observe that the impact of model error on DDTD is very mild. In some cases such as rank-3 DDTD in Maze, between $\theta=0.3$ and $\theta=0.5$, we observe slightly slower convergence with higher model error. The hyperparamaters of TD Learning and DDTD are given in Tables 1,2, and 3.
Figure 14: Comparison of DDTD with TD Learning and Dyna in Maze with $\gamma=0.99$ (Top) and $\gamma=0.95$ (Bottom). Left: low model error with $\theta=0.1$. Middle: medium model error with $\theta=0.3$. Right: high model error with $\theta=0.5$. Each curve is average of 20 runs. Shaded area shows one standard error.
Table 1: Hyperparamters for the Maze environment for $\gamma=0.99$. Cells with multiple values provide the value of the hyperparameter for $\theta=0.1$, $\theta=0.3$, and $\theta=0.5$, respectively.
@@
Table 2: Hyperparamters for the Maze environment for $\gamma=0.95$. Cells with multiple values provide the value of the hyperparameter for $\theta=0.1$, $\theta=0.3$, and $\theta=0.5$, respectively.
rank-1 DDTD rank-2 DDTD rank-3 DDTD rank-4 DDTD TD
$\eta$ (learning rate) $0.5,0.5,0.5$ $0.5,0.5,0.5$ $0.4,0.4,0.4$ $0.4,0.4,0.4$ 0.5
$\alpha$ $0.8,0.8,0.8$ $0.8,0.8,0.8$ $0.8,0.8,0.8$ $0.8,0.8,0.8$
$K$ 10 10 10 10 -
Figure 15: Comparison of DDTD with TD Learning and Dyna in Chainwalk with $\gamma=0.99$ (Top) and $\gamma=0.95$ (Bottom). Left: low model error with $\theta=0.1$. Middle: medium model error with $\theta=0.3$. Right: high model error with $\theta=0.5$. Each curve is average of 20 runs. Shaded area shows one standard error.
Table 3: Hyperparamters for the Chainwalk environment for both $\gamma=0.99$ and $\gamma=0.95$. Cells with multiple values provide the value of the hyperparameter for $\theta=0.1$, $\theta=0.3$, and $\theta=0.5$, respectively.